# UCINet0: A Machine Learning based Receiver for 5G NR PUCCH Format 0

### v1

@@ Abstract
Accurate decoding of Uplink Control Information (UCI) on the Physical Uplink Control Channel (PUCCH) is essential for enabling 5G wireless links. This paper explores an AI/ML-based receiver design for PUCCH Format 0. Format 0 signaling encodes the UCI content within the phase of a known base waveform and even supports multiplexing of up to 12 users within the same time-frequency resources. Our first-of-a-kind neural network classifier, which we term UCINet0, is capable of predicting when no user is transmitting on the PUCCH, as well as decoding the UCI content of any number of multiplexed users, up to 12. Inference results with both simulated and hardware-captured field datasets show that the UCINet0 model outperforms conventional DFT-based decoders across all SNR ranges.
Index Terms:
5G, AI/ML, Multi-label Classification, Neural Network, Physical Uplink Control Channel, Format 0
@@ I Introduction
The successful establishment of any wireless communication link between two entities relies on feedback signalling from both ends to indicate the channel quality as well as the status of previous transmissions. The Third Generation Partnership Project (3GPP) standards define the 5G NR (New Radio) Physical Uplink Control Channel (PUCCH) [1], which is the key enabler of such feedback in the uplink direction. It is a dedicated channel on which a User Equipment (UE) can send control information to a Base station (gNB). Uplink Control Information (UCI) carried by the PUCCH includes (1) Hybrid Automatic Repeat Request (HARQ) acknowledgements for prior downlink transmissions (Base station to UE), (2) Scheduling Request (SR) for the subsequent allocation of uplink transmission resources, and (3) Channel State Information (CSI) reports containing channel quality metrics that facilitate link adaptation, precoding, and downlink resource allocation.
In order to encompass the wide range of capacity and latency requirements of various 5G applications, the 5G standards [2, 3, 4] have provisioned five different formats of the PUCCH transmission (See Table I for a summary of the five formats). Smaller UCI payloads of 1 or 2 HARQ bits and/or an SR are transferred using Formats 0 and 1. Larger payloads, including several SR and HARQ bits, along with CSI reports, are transmitted using Formats 2, 3, and 4. Furthermore, Formats 0 and 2 are suitable for low latency applications since they occupy only 1 or 2 OFDM symbols in the time domain. Formats 1, 3, and 4 are suitable for applications requiring improved coverage and capacity since they extend between 4 and 14 OFDM symbols in the time domain and provide SNR gain.
@@ 4 HARQ, SR, CSI Long 4-14 1 DFT-S-OFDM $\frac{\pi}{2}$-BPSK or QPSK Yes Yes
I-A Our Contributions
In this paper, we focus on PUCCH Format 0, a particularly important format used during several steps of the 5G initial attach process [5]. Specifically, we look at the design and performance analysis of a Format 0 receiver using AI/ML techniques such as Neural Network classifiers. We make the following contributions.
  • PUCCH Format 0 uses a waveform-based technique for information transfer. It encodes the UCI within phase rotations applied to a pre-defined base sequence. This paper presents a Machine Learning based decoder for Format 0 that uses a Neural Network (NN) to predict the phase rotation. This paper, an extension of our previous work [6], is one of the first attempts at designing such a receiver for 5G PUCCH Format 0.
  • PUCCH Format 0 signals can be transmitted either by individual UEs or by multiple UEs (a maximum of 12 UEs [2]) multiplexed in the same time-frequency resources. Conventional decoding methods employ iterative correlation of the coherently combined signals with the known base sequences. This paper shows that a single NN model can replace the correlation operations, thus decoding the UCI from a single UE up to a maximum of 12 multiplexed UEs, with each UE transmitting any combination of HARQ and SR payloads.
  • One of the challenges of correlation-based methods is their susceptibility to false detections. Conventional approaches to determine whether a UE has actually transmitted on the PUCCH involve signal power-based thresholds, which are often determined either based on observation or tedious simulations. We show that the NN model described in this paper is robust enough to identify false transmissions, thus eliminating the need for thresholds.
  • This paper provides insights into the generation of training and testing wireless signal datasets derived from both MATLAB simulations and real-time over-the-air field captures taken at the indigenous 5G Testbed at IIT Madras [7].
@@ I-A Our Contributions
  • AI/ML models must adhere to certain complexity, memory, and latency constraints to become deployable on hardware devices such as Field Programmable Gate Arrays (FPGA). To this end, this paper provides a complexity analysis of the NN models presented.
@@ II Background on 5G NR PUCCH Format 0
PUCCH Format 0 is used to transfer one or two HARQ acknowledgments and/or an SR [4]. In 5G, Format 0 signalling is used as early as the initial attach procedure in which a UE is trying to latch onto a gNB. During the initial attach, there are several critical downlink transmissions, such as the Radio Resource Control (RRC) Setup, that need to be acknowledged, which the UE signals on PUCCH Format 0. It is to be noted that at this early stage of the connection process, neither the UE nor the gNB are expected to have obtained detailed information about the wireless surroundings. Consequently, Format 0 employs a sequence-based transmission in which phase-rotated versions of a pre-defined base sequence are transmitted by the UE. The UCI is encoded in the phase of the base sequence. Since the base sequence is assumed to be known at both ends, Format 0 signalling does not contain any pilot reference signals such as the Demodulation Reference Signals (DMRS), nor does it use Quadrature Amplitude Modulation (QAM) to modulate the UCI bits [8].
@@ II-A Time-Frequency Allocation, Sequence Generation and UCI Encoding
The cyclically shifted base sequence is mapped to the Resource Grid, where it occupies one Resource Block (RB) in the frequency domain and either one or two symbols in the time domain. One Resource Block in the frequency domain contains 12 Resource Elements (or 12 sub-carriers of the OFDM grid i.e., $N_{sc}^{RB}=12$). Since PUCCH Format 0 can span two symbols, either 12 or 24 Resource Elements can be occupied. The second symbol contains a sequence similar to that of the first symbol and can be used for SNR enhancement at the receiver. In more concrete terms, the PUCCH Format 0 sequence in the frequency domain includes a phase rotation $\alpha$ applied to a low Peak-to-Average Power Ratio (PAPR) base-sequence [2] $r_{u,v}^{\alpha}(k)$, and is given by,
@@
where $k=0,1,2,...,N_{sc}^{RB}-1$ and $\phi(k)$ is given by Table 5.2.2.2-2 in [2] which is a set of phase factors that lead to the generation of low PAPR sequences. The subscripts $u$ and $v$ represent the group number and the sequence number within the group, respectively. The values of $u$ and $v$ are configured by higher layer group hopping parameters. We note that when PUCCH Format 0 is two OFDM symbols long, intra-slot frequency hopping can be enabled. In this paper, for ease of exposition, we assume that intra-slot frequency hopping is disabled and utilize only one symbol for decoding the UCI bits.
@@
  • $m_{0}\in[0,11]$ is the initial cyclic shift that is used to differentiate various multiplexed users as defined in  [2].
@@
II-B PUCCH Format 0 Transmission Scenarios
While $m_{0}$ and $n_{cs}$ are drawn from higher layer (L2) configurations, the UE chooses the $m_{cs}$ depending on the UCI content it transfers. Consider the following two examples of possible PUCCH Format 0 transmission scenarios.
II-B1 Single UE Transmission
A single UE scheduled to transmit on a particular time-frequency allocation chooses an $m_{cs}$ value based on its UCI content as defined in Table. II. In the table, NACK and ACK refer to Negative and Positive Acknowledgments, respectively. +ve SR indicates an SR transmission and a -ve SR denotes that an SR is not transmitted despite the PUCCH resource allocation (the UE does not require uplink PUSCH allocations).
@@ 1 SR 0 (NACK, NACK, -ve SR),
II-B2 Multiplexed UE Transmission
In the interest of optimizing radio resources, PUCCH Format 0 has a unique potential to multiplex more than one UE on the same time-frequency locations. This is achieved by allocating different initial cyclic shifts ($m_{0}$) to each UE. Since the number of possible values of $m_{0}$ is limited to 12, the number of UEs that can be multiplexed is upper-bounded by 12. The maximum number of UEs that can be multiplexed is also driven by the UCI content. A few examples of this are shown in Figure 1. Figure 1a shows how assigning $m_{0}$ values of 0 through 5 to a maximum of 6 different UEs allows each of them to pick one of the two possible cyclic shifts and transmit 1 HARQ bit on the same resources. Similarly, Figure 1b shows that $m_{0}$ values of 0 through 2 allow for 3 UEs to be multiplexed when each of them transmits 2 HARQ bits (or 1 HARQ and 1 SR bit) with one of the four possible cyclic shifts. Additionally, multiplexing across different types of UCI content is possible, wherever applicable. For example, a UE transmitting 2 HARQ bits and an SR bit cannot be multiplexed with another transmitting the same combination but can be multiplexed with up to 3 more UEs transmitting SR bits or 1 more UE transmitting 2 HARQ bits (Figure 1c). Multiplexing of multiple UEs is possible as long as there are no overlaps in the assigned possible cyclic shifts for each of the UEs.
@@ II-B2 Multiplexed UE Transmission
Accounting for UE multiplexing, the generalized multi-user Format 0 received signal in the frequency domain is given by,
$y(k)=\sum_{m=0}^{N_{UE}-1}h_{m}(k)e^{j\alpha_{m}k}\cdot\bar{r}_{u,v}(k)+w(k),$ (3)
@@
where $k=0,1,2,...,N_{sc}^{RB}-1$, $N_{UE}$ is the number of multiplexed UEs in a given resource block, $h_{m}(k)$ is the channel between the gNB and the UE transmitting the base sequence with cyclic shift $\alpha_{m}$ in the $k^{th}$ Resource Element (observe that all multiplexed UEs use the same base sequence). The noise is denoted by $w(k)$.
II-C Existing Methods For PUCCH Format 0
II-C1 Conventional Approaches
@@
We reiterate that Format 0 has no provision for DMRS (pilots), and hence, there is no channel estimation or equalization possible. Blind correlation methods [9] [10] exist for such scenarios in which channel information is unknown, but these methods are often reserved for scenarios in which the resource allocation information is also not known. One example of such a scenario is the initial time synchronization between the gNB and the UE. Since the allocation of the synchronization sequences is unknown, longer correlations need to be performed. But, by the time the first Format 0 signal is received at the gNB, time and frequency synchronization are already achieved, and the gNB is aware of the UE’s transmission [11]. Since the allocation is known and is always 1 RB long, a simpler length 12 correlation is sufficient.
@@
In  [12, 13, 14], various receiver techniques based on correlation are considered. Correlation relies on the fact that the low PAPR base sequence used for encoding the UCI content is known at the receiver. The decoding method involves correlating the received samples with various cyclically shifted versions of the base sequence. The predicted cyclic shift is the cyclic shift that gives the highest correlation magnitude. It should be noted that peak selection, by definition, requires an optimal threshold for determining whether a peak is due to noise or a true transmission.
@@
The work in [12] shows that correlation-based signal detection outperforms methods based on raw signal power measurements even under fading channel scenarios. A normalized correlation-based solution for identifying false detections of HARQ bits is provided in [13]. The solution uses the probability distributions of a normalized correlation peak to determine the optimum threshold for classifying a reception as false. The PUCCH format 0 receiver algorithm proposed in [14] eliminates phase opposition across multiple hops by finding the sum of correlation magnitudes across symbols instead of correlation values.
In the 5G testbed deployment at IIT Madras [7], the correlation is implemented by an equivalent DFT-based algorithm. Later in the paper, we use this as a baseline for comparison. Such an algorithm proves to be better from a hardware perspective due to its optimized use of resources (avoiding the need to correlate with all the shifted base sequences), reduced latency, and higher throughput. In the IIT Madras 5G testbed, a DFT-based receiver for PUCCH Format 0 has been implemented on custom Field Programmable Gate Array (FPGA) boards for real-time operations.
Figure 2: Illustration of the output of Correlation of received Format 0 signal with the known base sequence for (a) 1 UE transmission and (b) 3 UEs multiplexed transmission
This method recovers the phase rotation $\alpha$ by taking the 12-point DFT of $y(n)\cdot\bar{r}^{*}_{u,v}(n)$. Since $y(n)=e^{j\alpha n}\cdot\bar{r}_{u,v}(n)$, the multiplication of the base sequence $\bar{r}_{u,v}$ with its complex conjugate forces it to unity. The 12-point DFT of the resultant exponential term $e^{j\alpha n}$ results in a peak at $\alpha$. This is shown in Figure 2(a) with a correlation peak at $index=4$. The correlation index corresponding to the maximum correlation amplitude is taken as the estimated $\alpha$ value. Recall from (2) that $\alpha$ is a combination of $m_{0}$, $m_{cs}$ and $n_{cs}$. Therefore, upon the subtraction of $m_{0}+n_{cs}$ from $\alpha$ (subtraction is modulo 12), the UCI-specific cyclic shift $m_{cs}$ remains. In scenarios where multiple users ($N_{UE}$ of them) transmit PUCCH Format 0 signals in the same time-frequency location, the DFT algorithm selects the top $N_{UE}$ peaks from the DFT output and uses them to determine the $m_{cs}$ for each of the $N_{UE}$ multiplexed UEs. Figure 2(b) shows an example in which 3 UEs are multiplexed. In this case, we select the top 3 peaks (at $index=4,6,$ and $8$) as $\alpha$ values.
We observe that while the above works in literature do provide false and missed detection performance analysis, they are based on fixed thresholds that are specific to a certain scenario and derived from lengthy Monte Carlo simulations. Another significant element that is not considered in [12, 13, 14] and considered only to some extent in [7] is the multiplexing of multiple UEs in the same resources.

### v2

@@ Abstract
Accurate decoding of Uplink Control Information (UCI) on the Physical Uplink Control Channel (PUCCH) is essential for enabling 5G wireless links. This paper explores an AI/ML-based receiver design for PUCCH Format 0. Format 0 signaling encodes the UCI content within the phase of a known base waveform and even supports multiplexing of up to 12 users within the same time-frequency resources. The proposed neural network classifier, which we term UCINet0, is capable of predicting when no user is transmitting on the PUCCH, as well as decoding the UCI content for any number of multiplexed users (up to 12). The test results with simulated, hardware-captured (lab) and field datasets show that the UCINet0 model outperforms conventional correlation-based decoders across all SNR ranges and multiple fading scenarios.
@@ I INTRODUCTION
The successful establishment of any wireless communication link between two entities relies on feedback signalling from both ends to indicate the channel quality as well as the status of previous transmissions. The Third Generation Partnership Project (3GPP) standards define the 5G New Radio (NR) Physical Uplink Control Channel (PUCCH) [1], which is the key enabler of such feedback in the uplink direction. It is a dedicated channel on which a User Equipment (UE) can send control information to a Base station (gNB). Uplink Control Information (UCI) carried by the PUCCH includes (1) Hybrid Automatic Repeat Request (HARQ) acknowledgements for prior downlink transmissions (gNB to UE), (2) Scheduling Request (SR) for the subsequent allocation of uplink transmission resources, and (3) Downlink Channel State Information (CSI) reports containing channel quality metrics that facilitate link adaptation, precoding, and downlink resource allocation.
In order to encompass the wide range of capacity and latency requirements of various 5G applications, the 5G standards [2, 3, 4] have provisioned five different formats of the PUCCH transmission. See Table I for a summary of the five different formats. Smaller UCI payloads of 1 or 2 HARQ bits and/or an SR are transferred using Formats 0 and 1. Larger payloads, including several SR and HARQ bits, along with CSI reports, are transmitted using Formats 2, 3, and 4. Furthermore, Formats 0 and 2 are suitable for low latency applications since they occupy only 1 or 2 OFDM symbols in the time domain. Formats 1, 3, and 4 are suitable for applications requiring improved coverage and capacity since they extend between 4 and 14 OFDM symbols in the time domain and can provide SNR gain.
@@ 4 HARQ, SR, CSI Long 4-14 1 DFT-S-OFDM $\frac{\pi}{2}$-BPSK or QPSK Yes Yes
In this paper, we focus on PUCCH Format 0, a particularly important format first used during several steps of the 5G initial attach process [1] and for subsequent feedback transmissions (small payload) even after the establishment of the communication link. Specifically, we look at the design and performance analysis of a receiver for Format 0 using AI/ML techniques such as Neural Network (NN) classifiers. In this paper, we make the following contributions:
  • PUCCH Format 0 uses a waveform-based technique for information transfer. It encodes the UCI within phase rotations applied to a pre-defined base sequence. This paper presents a Machine Learning based decoder for Format 0 that uses a neural network to predict the phase rotation. This paper, an extension of our previous work [5], is one of the early attempts at designing such a receiver for 5G PUCCH Format 0.
  • The integration of AI capabilities with communication systems can be done by replacing existing functionalities with AI-based approaches, by adding new AI-based functionalities, or by adding AI assistance to existing approaches. This paper proposes that the decoding of Format 0 could be performed by an AI/ML model, in lieu of conventional approaches such as correlation with the known base sequences. The AI/ML model takes only the extracted Format 0 resources from the Resource Grid as input and the functioning of this model is independent of the rest of the receiver chain.
  • Format 0 signals can be transmitted either by a single UE or by multiple UEs (a maximum of 12 UEs [2]) multiplexed in the same time-frequency resources. The AI/ML model described in this paper is capable of decoding the UCI from a single UE up to a maximum of 12 multiplexed UEs, with each UE transmitting any of the allowed combinations of HARQ and SR payloads.
  • One of the challenges of correlation-based methods is their susceptibility to false detections. Conventional approaches to determine whether a UE has actually transmitted on the PUCCH involve signal power-based thresholds, which are often determined either based on observation or tedious field measurements. We show that the UCINet0 model described in this paper is robust enough to identify false transmissions, thus eliminating the need for thresholds.
  • This paper provides insights into the generation of training and testing wireless signal datasets derived from both MATLAB simulations and real-time over-the-air captures taken at our indigenous 5G Testbed at IIT Madras [6].
@@
  • AI/ML models must adhere to certain complexity, memory, and latency constraints to become deployable on hardware devices such as Field Programmable Gate Arrays (FPGAs). To this end, this paper provides a complexity analysis of the NN models presented.
@@ II BACKGROUND ON 5G NR PUCCH FORMAT 0
PUCCH Format 0 is used to transfer one or two HARQ acknowledgments and/or an SR [4]. In 5G, the use of Format 0 signalling begins as early as the initial attach procedure in which a UE is trying to latch onto a gNB. During the initial attach, there are several critical downlink transmissions, such as the Radio Resource Control (RRC) Setup, that need to be acknowledged, which the UE signals using PUCCH Format 0. It is to be noted that at this early stage of the connection process, neither the UE nor the gNB is expected to have obtained detailed information about the wireless surroundings. Consequently, Format 0 employs a sequence-based transmission in which phase-rotated versions of a pre-defined base sequence are transmitted by the UE. The UCI is encoded in the phase of the base sequence. Since the base sequence is assumed to be known at both ends, Format 0 signalling does not contain any pilot reference signals such as the Demodulation Reference Signals (DMRS), nor does it use Quadrature Amplitude Modulation (QAM) to modulate the UCI bits [7].
@@ II-A TIME-FREQUENCY ALLOCATION, SEQUENCE GENERATION and UCI ENCODING
The cyclically shifted base sequence is mapped to the Resource Grid, where it occupies one Resource Block (RB) in the frequency domain and either one or two symbols in the time domain. Note that, one Resource Block in the frequency domain contains 12 Resource Elements (or 12 sub-carriers of the OFDM grid i.e., $N_{sc}^{RB}=12$). Since PUCCH Format 0 can span two symbols, either 12 or 24 Resource Elements can be occupied. The second symbol contains a sequence similar to that of the first symbol and can be used for SNR enhancement at the receiver. In more concrete terms, the PUCCH Format 0 sequence in the frequency domain includes a phase rotation $\alpha$ applied to a low Peak-to-Average Power Ratio (PAPR) base-sequence [2] $r_{u,v}^{\alpha}(k)$, and is given by,
@@
where $k=0,1,2,...,N_{sc}^{RB}-1$ and $\phi(k)$ is given by Table 5.2.2.2-2 in [2], which is a set of phase factors that lead to the generation of low PAPR sequences. The subscripts $u$ and $v$ represent the group number and the sequence number within the group, respectively. The values of $u$ and $v$ are configured by higher layer group hopping parameters. We note that when PUCCH Format 0 is two OFDM symbols long, intra-slot frequency hopping can be enabled. In this paper, for ease of exposition, we assume that intra-slot frequency hopping is disabled and utilize only one symbol for decoding the UCI bits.
@@
  • $m_{0}\in\{0,1,2,\dots,11\}$ is the initial cyclic shift that is used to differentiate various multiplexed users as defined in  [2].
@@
While $m_{0}$ and $n_{cs}$ are drawn from higher layer (L2) configurations, a UE scheduled to transmit on a particular time-frequency allocation chooses the $m_{cs}$ depending on the UCI content it transfers, as defined in the Table. II. In the table, NACK and ACK refer to Negative and Positive Acknowledgments, respectively. Positive SR (i.e., +ve SR) indicates an SR transmission, and a Negative SR (i.e., -ve SR) denotes that an SR is not transmitted despite the PUCCH resource allocation (the UE does not require Physical Uplink Shared Channel (PUSCH) allocations).
@@ 1 SR 0 (NACK, NACK, -ve SR),
II-B PUCCH FORMAT 0 TRANSMISSION SCENARIOS
In the interest of optimizing radio resources, PUCCH Format 0 has a unique potential to multiplex more than one UE on the same time-frequency locations. This is achieved by allocating different initial cyclic shifts ($m_{0}$) to each UE. Since the number of possible values of $m_{0}$ is limited to 12, the number of UEs that can be multiplexed is upper-bounded by 12. The maximum number of UEs that can be multiplexed is also driven by the UCI content. A few examples of this are shown in Figure 1. Figure 1a shows how assigning $m_{0}$ values of 0 through 5 to a maximum of 6 different UEs allows each of them to pick one of the two possible cyclic shifts and transmit 1 HARQ bit on the same resources. Similarly, Figure 1b shows that $m_{0}$ values of 0 through 2 allow for 3 UEs to be multiplexed when each of them transmits 2 HARQ bits (or 1 HARQ and 1 SR bit) with one of the four possible cyclic shifts. Additionally, multiplexing across different types of UCI content is possible, wherever applicable. For example, a UE transmitting 2 HARQ bits and an SR bit cannot be multiplexed with another transmitting the same combination, but can be multiplexed with up to 4 more UEs transmitting SR bits or 1 more UE transmitting 2 HARQ bits (Figure 1c). Multiplexing of multiple UEs is possible as long as there are no overlaps among the possible cyclic shifts assigned for each UE.
@@ II-C MATHEMATICAL MODEL OF PUCCH FORMAT 0
Accounting for UE multiplexing, the generalized equation for the received multi-user Format 0 signal, in the frequency domain is given by,
$y(k)=\sum_{m=0}^{N_{UE}-1}\beta_{m}h_{m}(k)e^{j\alpha_{m}k}\bar{r}_{u,v}(k)+w(k),$ (3)
where $k=0,1,2,...,N_{sc}^{RB}-1$. Here, $\beta_{m}$ indicates whether the $m$-th UE, which was scheduled to transmit on the PUCCH Format 0, transmitted or not, i.e., $\beta_{m}\in\{0,1\}$. It should be observed that while $N_{UE}$ UEs are scheduled, all of the UEs need not transmit and hence the actual number of transmissions $\tilde{N_{UE}}$ can be less than or equal to $N_{UE}$, i.e.,
$\tilde{N}_{UE}=\sum_{m=0}^{N_{UE}-1}\beta_{m}\leq N_{UE}.$ (4)
Let
$\Delta=N_{UE}-\tilde{N}_{UE}.$ (5)
@@
In Eq. 3, $h_{m}(k)$ is the channel between the gNB and the UE transmitting the base sequence with cyclic shift $\alpha_{m}$ in the $k^{th}$ Resource Element (observe that all multiplexed UEs use the same base sequence). The noise is denoted by $w(k)$.
The goal of PUCCH Format 0 decoding is to determine $\alpha_{m}$ ($m=0,1,2,\dots,N_{UE}-1$), given that the gNB observes $y(k)$ ($k=0,1,2,...,N_{sc}^{RB}-1$). The corresponding optimization problem is shown in the equations below.
[…]
$\sum_{m=0}^{N_{UE}-1}\beta_{m}\leq N_{UE},$
$\overline{\alpha}\in A_{0}\times A_{1}\times\ldots\times A_{N_{UE}-1}.$
Here, $A_{i}=\mathcal{R}(\overline{C}_{i},{m_{0}}_{i},n_{cs})$ is a set containing $\alpha_{i}$ values for user $i$, given its UCI content vector $\overline{C}_{i}$, its initial cyclic shift $m_{0_{i}}$ and $n_{cs}$ values. Note that $i=0,1,\dots,N_{UE}-1$. For $N_{UE}$ multiplexed users, given their UCI content vectors, their initial cyclic shift and $n_{cs}$ values, the set $\overline{\alpha}$ contains all possible combinations of cyclic shifts $\alpha_{i}$ ($i=0,1,\dots,N_{UE}-1$). These combinations can be represented as $N_{UE}$-length tuples and the set of all combinations $\overline{\alpha}$ can be obtained by the Cartesian Product of the sets $A_{i}$ where $i=0,1,\dots,N_{UE}-1$.
Consider the following illustrative examples.
Example 1: Assume that $N_{UE}=1$, $n_{cs}=0$, $m_{0_{0}}=0$ and the UCI content for the scheduled UE is 1 SR and 1 HARQ i.e., $\overline{C}_{0}=[1,1]$. Then,
$\overline{\alpha}\in A_{0}=\mathcal{R}([1,1],0,0)=\{0,3,6,9\}.$
Example 2: Assume that $N_{UE}=2$, $n_{cs}=0$ for both UEs, $m_{0_{0}}=0$, $m_{0_{1}}=1$ and the UCI content for the first UE is 1 SR and 1 HARQ i.e., $\overline{C}_{0}=[1,1]$ and that for the second UE is 1 HARQ i.e., $\overline{C}_{1}=[0,1]$. Then,
$A_{0}=\mathcal{R}([1,1],0,0)=\{0,3,6,9\},$
$A_{1}=\mathcal{R}([0,1],1,0)=\{1,7\},$
$\overline{\alpha}\in A_{0}\times A_{1}=\{(0,1),(0,7),(3,1),$
$(3,7),(6,1),(6,7),(9,1),(9,7)\}.$
Example 3: Assume that $N_{UE}=3$, $n_{cs}=0$ for all the three UEs, $m_{0_{0}}=0$, $m_{0_{1}}=1$, $m_{0_{2}}=2$ and the UCI content for the first UE is 1 SR and 1 HARQ i.e., $\overline{C}_{0}=[1,1]$, that for the second UE is 1 HARQ i.e., $\overline{C}_{1}=[0,1]$ and that for the third UE is 1 SR i.e., $\overline{C}_{2}=[1,0]$. Then,
$A_{0}=\mathcal{R}([1,1],0,0)=\{0,3,6,9\},$
$A_{1}=\mathcal{R}([0,1],1,0)=\{1,7\},$
$A_{2}=\mathcal{R}([1,0],2,0)=\{2\},$
$\overline{\alpha}\in A_{0}\times A_{1}\times A_{2}=\{(0,1,2),(0,7,2),(3,1,2),$
$(3,7,2),(6,1,2),(6,7,2),(9,1,2),(9,7,2)\}.$
II-D CONVENTIONAL RECEIVER METHODS FOR PUCCH FORMAT 0
@@
Since Format 0 PUCCH has no provision for pilots, coherent detection with channel estimation is not possible. Blind correlation methods [8] [9] exist for scenarios in which channel information is unknown, but these methods are often reserved for scenarios in which resource allocation information is also not known. An example of such a scenario is the initial time synchronization between the gNB and the UE. Since the allocation of the synchronization sequences is unknown, longer correlations need to be performed. However, by the time the first Format 0 signal is received at the gNB, time and frequency synchronization is already achieved, and the gNB is aware of the UE’s transmission [10]. Since the allocation is known and is always 1 RB long, a simpler length-12 correlation is sufficient.
@@
In  [11, 12, 13], various receiver techniques based on correlation are considered. The correlation relies on the fact that the low PAPR base sequence used to encode the UCI content is known to the receiver. The decoding method involves correlating the received samples with various versions of the base sequence that are cyclically shifted. The predicted cyclic shift is the cyclic shift that gives the highest correlation magnitude. It should be noted that peak selection, by definition, requires an optimal threshold to determine whether a peak is due to noise or a true transmission.
@@
The work in [11] shows that correlation-based signal detection outperforms methods based on raw signal power measurements even in fading channel scenarios. A normalized correlation-based solution for identifying false detections of HARQ bits is provided in [12]. The solution uses the probability distributions of a normalized correlation peak to determine the optimum threshold to classify a reception as false. The receiver algorithm for PUCCH Format 0 proposed in [13] eliminates phase opposition across multiple hops by finding the sum of correlation magnitudes across symbols instead of correlation values.